OpenStudy (anonymous):

the part of the parabola y=4-x^2 from x=0 to x=2 is rotated about the y axis to form a surface. A cone is inscribed in the resulting paraboloid with its vertex at the origin and its base touching the parabola. At what radius and altitude does the max volume occur? What is the max volume?

OpenStudy (anonymous):

What is your question?

OpenStudy (anonymous):

the part of the parabola y=4-x^2 from x=0 to x=2 is rotated about the y axis to form a surface. A cone is inscribed in the resulting paraboloid with its vertex at the origin and its base touching the parabola. At what radius and altitude does the max volume occur? What is the max volume?

OpenStudy (anonymous):

Yes that makes sense. Thank you so much for your help!

OpenStudy (anonymous):

Wait, There is a small correction.

OpenStudy (anonymous):

OK. Thanks.

OpenStudy (anonymous):

It does not change the answer. I will rewrite the new version.

OpenStudy (anonymous):

\[ V(x) =\frac 1 3 x(4-x^2)=\frac{4 x}{3}-\frac{x^3}{3}\\ V'(x)=\frac{4}{3}-x^2\\ V''(x)=-3 x\\ V'(x) = 0, \implies x=\pm \frac 2 {\sqrt 3} \]

OpenStudy (anonymous):

We take the positive value of x and at that point the second derivative is negative so we have a maximum

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

YW